the midpoint of AB is M (3, 2). if the coordintes of A are (8, 7). what are the coordintes
Answer by greenestamps(13334)



Incomplete post; presumably you want to find the coordinates of point B.

Instead of using formal algebra, imagine walking along AB from A towards B. M is the midpoint of AB, which is the halfway point of your "walk".

From A to your halfway point M, you have moved from 8 to 3 in the x direction, a change of -5. To get to the end of your walk, you need to go another -5 units in the x direction; that puts you at x = 3-5 = -2.

From A to your halfway point M, you have moved from 7 to 2 in the y direction, also a change of -5. To get to the end of your walk, you need to go another -5 units in the y direction; that puts you at y = 2-5 = -3.

At the end of your walk, you are at (x,y) = (-2,-3).

ANSWER: the coordinates of B are (-2,-3)

If your assignment is to find the answer using formal algebra....

Let the coordinates of B be (x,y). Then according to the midpoint formula....

(x+8)/2 = 3
x+8 = 6
x = 6-8 = -2

and

(y+7)/2 = 2
y+7 = 4
y = 4-7 = -3

Again (of course), point B is (-2,-3)